### v2

# VTG-LLM: Integrating Timestamp Knowledge into Video LLMs for Enhanced Video Temporal Grounding

Abstract
Video Temporal Grounding (VTG) focuses on accurately identifying event timestamps within a particular video based on a linguistic query, playing a vital role in downstream tasks such as video browsing and editing. While Video Large Language Models (video LLMs) have made significant progress in understanding video content, they often face challenges in accurately pinpointing timestamps within videos, which limits their performance on VTG tasks. Therefore, to improve video LLMs’ ability to effectively locate timestamps, we argue that two critical aspects need to be enhanced. First, it is essential to have high-quality instructional tuning datasets that encompass mainstream VTG tasks. Second, directly incorporating timestamp knowledge into video LLMs is crucial, as it enables models to efficiently comprehend timestamp information. To address these needs, we first introduce VTG-IT-120K, a high-quality and comprehensive instruction tuning dataset that covers VTG tasks such as moment retrieval, dense video captioning, video summarization, and video highlight detection. Furthermore, we propose a specially designed video LLM model for VTG tasks, VTG-LLM, which (1) effectively integrates timestamp knowledge into visual tokens; (2) incorporates absolute-time tokens that specifically handle timestamp knowledge, thereby avoiding concept shifts; and (3) introduces a lightweight, high-performance slot-based token compression method to facilitate the sampling of more video frames. Comprehensive experiments showcase the superior performance of VTG-LLM in comparison to other video LLM methods across various VTG tasks. Our code and datasets are available at https://github.com/gyxxyg/VTG-LLM.
1 Introduction
Large language models (LLMs) (Kaplan et al., 2020; Achiam et al., 2023; Touvron et al., 2023) have exhibited considerable potential in capturing knowledge and tackling real-world challenges using a zero-shot approach. Recently, research has explored integrating knowledge from other modalities, such as vision (Liu et al., 2024; Li et al., 2023a) and audio (Ghosal et al., 2023), to enhance the capabilities of LLMs. Within the visual modality, video large language models (video LLMs) have emerged as a significant research area (Lin et al., 2023a; Zhang et al., 2023; Li et al., 2023b). These models aim to generate textual outputs based on given videos and text prompts, achieving remarkable success in multiple video-related tasks like video captioning (Xu et al., 2024) and question-answering (Li et al., 2023c).
Video Temporal Grounding (VTG) is a crucial component of video understanding. It requires models to accurately pinpoint event timestamps within a video based on the given query. This task is essential for subsequent operations such as video browsing and editing. The sub-tasks within VTG can be categorized based on output formats, as described by (Lin et al., 2023b). For instance, moment retrieval and dense video captioning (Fabian Caba Heilbron & Niebles, 2015; Zhou et al., 2018) tasks demand models to generate one or multiple timestamp intervals accompanied by captions; video highlight detection (Lei et al., 2021) task requires models to produce a salient score curve; and video summarization (Song et al., 2015) task calls for models to output a series of timestamps along with their corresponding video frames.
Despite the remarkable achievements of video LLMs in video understanding tasks, current models struggle to accurately pinpoint timestamps, leading to subpar performance on VTG tasks. Therefore, it is essential to enhance video LLMs by equipping them with the ability to effectively comprehend timestamps. To accomplish this, we contend that two key aspects need to be addressed. On one hand, high-quality instructional tuning datasets specifically tailored for VTG tasks are necessary. Existing datasets are either of low quality (Zellers et al., 2021) or exhibit significant task imbalance (Ren et al., 2023; Wang et al., 2024b). On the other hand, it is crucial to develop new models that efficiently and explicitly incorporate timestamp knowledge into video LLMs.
In this paper, we tackle the aforementioned issues by first introducing a high-quality instruction tuning dataset, VTG-IT-120K, which consists of 47.2K videos and 120K annotations. This dataset encompasses most mainstream VTG tasks, including moment retrieval (63.2K), dense video captioning (37.2K), video summarization (15.2K), and video highlight detection (3.9K). The VTG-IT-120K dataset is more task-balanced compared to existing datasets (Ren et al., 2023; Wang et al., 2024b), which either contain annotations exclusively for the moment retrieval task or provide a mere 75 samples for the video summarization task. Moreover, the VTG-IT-120K includes 51.9K self-collected data annotations of superior quality. Furthermore, we present a novel video LLM model, VTG-LLM, which comprises three components for efficiently integrating timestamp knowledge into video LLMs: (1) To ensure that the inputs of LLMs contain sufficient timestamp information, we introduce a sequence-time embedding method that integrates timestamp data into visual tokens; (2) Concept shifts can occur when different input data lead to the same output targets, which can obscure the decision boundary (Moreno-Torres et al., 2012). For instance, the number 20 could represent seconds or number of people. To mitigate this, we incorporate absolute-time tokens that specifically handle timestamp knowledge, without introducing quantization errors. This ensures that the LLM tokenizer and LLM prediction head do not confuse timestamp-related information with other digit-related information, such as counting and mathematics; (3) Given the limited context length of LLMs, the number of visual tokens must be reduced to allow for more video frame sampling. As such, we employ a lightweight slot-based token compression method to decrease the number of visual tokens to a fixed length, enabling video LLMs to sample more video frames. Numerical results illustrate the superior performance of the VTG-LLM compared to state-of-the-art video LLM methods. Additionally, ablation studies reveal that absolute-time tokens and sequence-time embedding significantly enhance the model’s performance in accurately locating timestamps. Furthermore, the slot-based token compression method outperforms the naive token compression baselines. Our key contributions are summarized as follows:
Figure 1: Overview of the VTG-LLM model. We utilize a fixed visual encoder and a Qformer model to encode sampled video frames into visual tokens. Subsequently, the absolute time $t_{i}$ and sequence ID $f_{i}$ are embedded and added to the top of each visual token for frame $i$. The slot-based token compression is designed to compress the number of tokens to a fixed number of visual tokens. To prevent overfitting and facilitate training, we employ a variety of prompts and well-structured answers. The timestamps with green color are encoded using six distinct time tokens.
  • We introduce VTG-IT-120K, a high-quality instruction tuning dataset comprising annotations for four distinct VTG tasks: moment retrieval (63.2K), dense video captioning (37.2K), video summarization (15.2K), and video highlight detection (3.9K).
  • We present a carefully crafted model, VTG-LLM, which incorporates timestamp information into video LLMs, providing models with the ability to understand timestamps. Quantitative results demonstrate the superior performance of VTG-LLM across various datasets, such as Charades-STA, QVHighlights, and YouCook2. Moreover, all three components are necessary for VTG-LLM to achieve the best results.
2 Related Works
Video Temporal Grounding (VTG) aims to accurately locate the timestamps of events within a given video (Lin et al., 2023b). This encompasses tasks such as moment retrieval (Gao et al., 2017; Zala et al., 2023b; Oncescu et al., 2021; Hendricks et al., 2018a), dense video captioning (Zellers et al., 2021; Zala et al., 2023b; Tang et al., 2019; Fabian Caba Heilbron & Niebles, 2015), video summarization (Song et al., 2015; Gygli et al., 2014), and video highlight detection (Lei et al., 2021; Xiao et al., 2023). Traditional methods mainly address VTG tasks through large-scale video-text pre-training, using training objectives such as video-text contrastive learning (Xu et al., 2021; Wang et al., 2022; Yan et al., 2022), video-text matching (Li et al., 2023d; Chen et al., 2024), and masked auto-encoding (Tong et al., 2022; Zhao et al., 2024). Although these methods have shown satisfactory results, they require resource-consuming pre-training, lack zero-shot capabilities, and often need further fine-tuning on many downstream tasks.
Recently, Video Large Language Models (video LLMs) have shown promise in video understanding tasks. Traditional video LLMs mainly generate captions that summarize videos (Zhang et al., 2023; Lin et al., 2023a; Li et al., 2023b; Maaz et al., 2023; Zhu et al., 2023). However, these models lack the capacity to accurately locate the timestamps of events within a given video. Some studies have attempted to address this issue. For example, TimeChat (Ren et al., 2023) constructs time-sensitive instruction tuning datasets and encodes timestamp knowledge into visual tokens. VTimeLLM (Huang et al., 2023) proposes a LLaVA-like three-stage training method. LITA (Huang et al., 2024) introduces fast-slow visual tokens and adds time tokens to LLM tokenizers. Momentor (Qian et al., 2024) proposes a time encoder to solve the quantization error of time tokens. HawkEye (Wang et al., 2024b) constructs a high-quality instruction tuning dataset based on InternVid (Wang et al., 2022) for the moment retrieval task. However, existing instruction tuning datasets for VTG tasks are either low-quality or exhibit extreme task imbalance, hindering model performance on overall tasks. Therefore, we propose a high-quality dataset VTG-IT to address this issue. Furthermore, we introduce VTG-LLM, which contains three well-designed components to boost video LLM performance on VTG tasks.
3 Method
In this section, we introduce the VTG-LLM model and the VTG-IT-120K dataset. The overview of the VTG-LLM model can be found in Figure 1.
3.1 VTG-LLM: Incorporating Knowledge about Timestamps into Video LLMs
In this subsection, we introduce the detailed structure of VTG-LLM, which includes three key components designed to help video LLMs better understand timestamps. Specifically, we propose: (1) a sequence-time embedding mechanism to straight-forwardly incorporate timestamp information into visual tokens; (2) the introduction of absolute-time token without quantization error to distinguish time-related knowledge from other digit-related knowledge; and (3) the use of slot-based token compression to enable video LLMs to effectively handle more frames. We will introduce these components in detail in the following parts of this section.
3.1.1 Sequence-Time Embedding
Sequence embedding.
Existing studies (Zhang et al., 2023; Ren et al., 2023) employ sequence embedding to incorporate relative time information into visual tokens. Specially, given $N\times M$ tokens $\{\mathbf{z}_{i,j}|1\leq i\leq N,1\leq j\leq M\}$ sampled from $N$ frames, the process of adding sequence embedding onto visual tokens can then be represented by:
$\hat{\mathbf{z}}_{i,j}=\mathbf{z}_{i,j}+[\mathbf{W}_{s}]_{i}\,,$ (1)
where $\mathbf{W}_{s}\in\mathbb{R}^{N\times d}$ is the weight of sequence-embedding matrix.
Sequence-Time Embedding.
Although promising, the sequence embedding only contains information about the temporal order and may not accurately represent knowledge about the absolute timestamps. For instance, sampled frames may not be uniformly distributed throughout the entire video (Ren et al., 2023; Huang et al., 2023), and the start time of the first sampled frame is also unknown. This makes it difficult to infer the timestamps of the frames. To address this issue, in addition to the sequence embedding, we also add the absolute time embedding as shown in the following equation
$\hat{\mathbf{z}}_{i,j}=\mathbf{z}_{i,j}+[\mathbf{W}_{s}]_{i}+[\mathbf{W}_{t}]_{t}\,.$ (2)
Here, $t$ is the absolute timestamp (in seconds) of the corresponding frames. $\mathbf{W}_{t}\in\mathbb{R}^{T\times d}$ is the absolute time embedding. $T$ is the maximum timestamp (in seconds), which is set to 8192 by default, supporting videos for more than 1 hours long theoretically. It’s important to note that to avoid disrupting the visual tokens generated by pre-trained visual encoders and to speed up the convergence, $\mathbf{W}_{t}$ is initialized by setting all entries to zero. Without the zero-initialization method, the performance of VTG-LLM significantly degraded, as shown in the "TE Random Initialize" part of Table 2.
Due to the imbalanced nature of video lengths and the gap between training and test data, there may be timestamps for which the absolute time embedding has not been trained. To address this issue, we further introduce a test-time interpolation mechanism. In detail, for the timestamp $t$ which is not trained during training, we first find the timestamps $t_{l}$ and $t_{r}$ that satisfy:
$t_{l}=\operatorname*{arg\,max}_{t_{l}}t_{1}<t,t_{l}\in\mathcal{T}_{tr}\,,t_{r}=\operatorname*{arg\,min}_{t_{r}}t_{r}>t,t_{r}\in\mathcal{T}_{tr}\,,$ (3)
where $\mathcal{T}_{tr}$ is the set of timestamps with trained absolute-time embedding. Then the absolute-time embedding of timestamp $t$ is given by
$[\mathbf{W}_{t}]_{t}=\frac{t-t_{l}}{(t_{r}-t_{l})}[\mathbf{W}_{t}]_{t_{l}}+\frac{t_{r}-t}{(t_{r}-t_{l})}[\mathbf{W}_{t}]_{t_{r}}\,.$ (4)
Remark 3.1 (Discussion on existing techniques that integrating information of timestamps into visual tokens.).
In addition to incorporating time embeddings into visual tokens, there are other techniques for integrating timestamp information into visual tokens. For instance, adding text inputs into Qformer (Ren et al., 2023) and inserting time tokens before visual tokens (Hua et al., 2024). However, we believe that these techniques are orthogonal to the time embedding approach, and it is also possible to combine these approaches in future work.
3.1.2 Absolute-Time Tokens
The integration of special time tokens into the tokenizer has proven advantageous for video temporal grounding tasks, as demonstrated in several studies (Yang et al., 2023; Qian et al., 2024; Huang et al., 2024). However, the use of relative time tokens (frame ID) has revealed certain limitations. Firstly, there is the issue of quantization errors, which prevent the models from predicting timestamps outside the sampled frames. Secondly, the token embedding requires training from scratch, and any changes to the number of time tokens necessitate further training. As a result, our goal is to develop a new time token mechanism that can both represent fine-grained time without quantization errors and represent a broad range of timestamps without requiring additional effort when sampling varying numbers of frames.
Absolute-time tokens for resolving the quantization errors.
To resolve the issue of quantization errors, we introduce the concept of absolute-time tokens. As depicted in Figure 1, we have incorporated eleven time tokens into the tokenizer. These consist of ten digit time tokens representing the digits from 0 to 9, and an additional token for a decimal point. All timestamps are represented using six time tokens. For instance, the time 120.5 seconds would be formatted as <TIME_ZERO><TIME_ONE><TIME_TWO><TIME_ZERO><TIME_DOT><TIME_FIVE>. This formulation allows the time tokens to handle videos more than 1 hours in length, with the precision remaining constant regardless of video length increases. Importantly, it is essential to format all timestamps using the same number of time tokens. As shown in the "Time Token not Formatted" section of Table 2, we found that maintaining a consistent format for all timestamps significantly improves model performance.
Initialization of token embedding for absolute-time tokens.
Although using the absolute time token resolves the quantization error, we discovered that randomly initialized time tokens negatively impact the original token embedding space, preventing the LLMs from learning accurate knowledge about the time tokens. This resulting in poor performance of models, as shown in the "Token Embedding not Initialized" part of Table 2. To tackle this issue, our goal is to initialize the time-related knowledge using the digit-related knowledge, as the digit-related knowledge has already been well-trained during the LLM pretraining stage.
This is achieved by adjusting the initialization of the weights of token embedding and LLM prediction head. Specifically, we consider the token embedding matrix $\mathbf{W}_{token}$. When initializing the embedding for the time token "<TIME_ONE>", we align the embedding of this time token with the embedding of the token "1":
$[\mathbf{W}_{token}]_{\text{ID}(\text{'<TIME\_ONE>'}),j}=[\mathbf{W}_{token}]_{\text{ID}(\text{'1'}),j}\,,\forall 0\leq j<d\,,$ (5)
where ID() denotes the token ID of the given token string. Similarly, for the LM prediction head, we employ the same method to transfer the knowledge from the digit tokens to the time tokens.
3.1.3 Slot-Based Token Compression
Compressing the number of visual tokens is a critical issue in video LLMs. On one hand, using all the visual tokens generated by visual backbones preserves all the information. However, due to the restrictions on LLM context length, the number of sampled frames must also be reduced. On the other hand, sampling a subset of frames may result in overlooking important information. To resolve this issue, we propose a simple yet effective method, namely slot-based token compression, to compress the visual tokens to a fixed number of tokens.
(a) Overview of VTG-IT
(b) Slot-Based Token Compression
Figure 2: Overview of VTG-IT and slot based sampler. As shown in Figure 2(a), the VTG-IT contains annotations of 4 important video temporal grounding tasks. The VTG-IT-X indicate self-collected annotations. Figure 2(b) shows the architecture of slot-based token compression. The $N$ tokens are passed through a trainable slot dispatcher, which assigns dispatch weights $\bm{\omega}\in\mathbb{R}^{N}$ to each slot. The tokens are then mixed based on these dispatch weights to obtain the slots.
As illustrated in Figure 2(b), given $N$ visual tokens $\mathbf{z}_{1},\cdots,\mathbf{z}_{N}$, and the trainable slot dispatcher $\mathbf{\Phi}\in\mathbb{R}^{K\times d}$, where $K$ represents the number of slots, the visual tokens first pass through the slot dispatcher to obtain the dispatch weights, and the tokens are mixed based on the dispatch weights to generate the slots. For instance, the slot $k$ is generated by
$\mathbf{s}_{k}=\sum_{i=1}^{N}\frac{\exp(\mathbf{\Phi}_{k}^{T}\mathbf{z}_{i})}{\sum_{j=1}^{N}\exp(\mathbf{\Phi}_{k}^{T}\mathbf{z}_{j})}\mathbf{z}_{i}\,.$ (6)
The slots are subsequently fed into the visual projection layers and serve as visual tokens.
Remark 3.2 (Discussion on slot mechanism).
The slot-based token compression is inspired by the slot mechanism in SoftMoE (Puigcerver et al., 2023). However, in SoftMoE, the slots are not used to reduce sequence length, which is distinct from our primary objective. Moreover, some existing multi-modal LLMs compress the token number through cross-attention (Zhang et al., 2023; Bai et al., 2023; Ren et al., 2023). Our approach involves training only one matrix $\mathbf{\Phi}$, making it more computationally efficient and less data-consuming, and achieve better performance (Table 2).
3.2 VTG-IT-120K: Formatted Time-Sensitive Instruction Tuning Dataset
In this section, we introduce the VTG-IT-120K, which consists of 51.9K self-collected annotations and 67.6K annotations obtained from TimeIT (Ren et al., 2023). As shown in Figure 2(a), VTG-IT-120K primarily covers four distinct video temporal grounding tasks, including:
  • Moment Retrieval (63.2K): For the moment retrieval task, we use HiRESTgrounding (Zala et al., 2023a), QuerYD (Oncescu et al., 2021), DiDeMo (Hendricks et al., 2018b), and self-collected VTG-IT-MR.
  • Dense Video Captioning (37.2K): For the dense video captioning task, we use HiRESTstep (Zala et al., 2023a), COIN (Tang et al., 2019), ActivityNet Captions (Fabian Caba Heilbron & Niebles, 2015), and self-collected VTG-IT-DVC.
  • Video Summarization (15.2K): For the video summarization task, we use TVSum (Song et al., 2015), SumMe (Gygli et al., 2014), and self-collected VTG-IT-VS.
  • Video Highlight Detection (3.9K): For the video highlight detection task, we use self-collected VTG-IT-VHD.
The VTG-IT-X annotations for the four tasks are self-collected using 16K videos from the YT-Temporal-180M dataset (Zellers et al., 2021).
Data formatting.
All the tasks are formatted to QA pairs, and we use diverse questions to help models understand human intention. The answers are formatted for better transferring knowledge between different tasks. In detail, we reformat the annotations of different tasks, and use more similar answer formats among tasks. More detailed examples can be found in Appendix B.
  • Moment retrieval and dense video captioning. The GPT responses for these two tasks are formatted as ’start time - end time, description’. The moment retrieval task will consist of a single sentence, while the dense video captioning task may include multiple sentences.
  • Video summarization and video highlight detection. The responses for these two tasks are formatted as ’timestamp, score, description’. For the video highlight detection task, the description is the same as the query; for the video summarization task, the descriptions use placeholders like ’the highlight of the video’.
4 Experiments
4.1 Experiment Settings
Models and training configuration.
Unless otherwise specified, we employ ViT-G/14 from EVA-CLIP (Sun et al., 2023) and the Qformer from InstructBLIP (Dai et al., 2024) as the visual backbone. The vision projection layer and the weights of sequence embedding are initialized using the Video-LLaMA checkpoint (Zhang et al., 2023). For the language model backbone, we utilize LLaMA-2-7B (Touvron et al., 2023). The number of slots is set to 256. The models are trained on the VTG-IT-120k and a randomly sampled subset (97k) from the Valley dataset (Luo et al., 2023b). We randomly sample 96 frames for each video. Training is carried out using 16 ATN 910B NPUs with a batch size of 64, taking approximately 40 hours to complete. The fine-tuning also uses 16 NPUs and set the batch size to 16. We set the learning rate to 3e-5 initially, and train the models for 10 epochs by default.
Evaluation datasets, metrics, and baseline models.
We evaluate the model performance on three different tasks:
  • dense video captioning. We employ Youcook2 (Zhou et al., 2018) as the test dataset, and following the evaluation settings in TimeChat (Ren et al., 2023). Metrics including CIDEr (Vedantam et al., 2015), which assessing the quality of the captions; (2) SODA_c (Fujita et al., 2020) for story-level evaluation; (3) F1 score to measure the model’s ability to accurately locate timestamps;
  • Moment retrieval. We utilize test set of Charades-STA (Gao et al., 2017) for the moment retrieval task and report the recall at IOU thresholds of 0.5 and 0.7.
  • Video highlight detection. We employ the validation set of the QVHighlights dataset (Lei et al., 2021) and report the mean average precision (mAP) with IOU thresholds of 0.5 and 0.75, as well as the HIT@1, which represents the hit ratio of the highest scored clip.
For baseline models, we select Valley (Luo et al., 2023b), VideoChat (Li et al., 2023b), Video-ChatGPT (Maaz et al., 2023), and Video-LLaMA (Zhang et al., 2023) as examples of traditional video LLMs. For video LLMs specifically designed for VTG tasks, we choose TimeChat (Ren et al., 2023), VTimeLLM (Huang et al., 2023), Momentor (Qian et al., 2024), and HawkEye (Wang et al., 2024b)  [^1] Although related, LITA (Huang et al., 2024) employs self-built evaluation sets, making fair comparison difficult..
4.2 Results
Zero-shot performance of the VTG-LLM on various VTG tasks.
In Table 1, we evaluate the performance of the VTG-LLM on three VTG tasks, including dense video captioning (Youcook2), moment retrieval (Charades-STA), and video highlight detection (QVHighlights). The results indicate that: (1) The VTG-LLM model outperforms all existing 7B video LLM models on the zero-shot setting across all three datasets; (2) The performance of the VTG-LLM on the Charades-STA dataset for IOU=0.7 surpasses that of the VTimeLLM-13B model, indicating the superior performance of the VTG-LLM in accurately locating timestamps. (3) The improvement of the VTG-LLM on metrics that evaluate timestamp location accuracy, such as the F1 score for the Youcook2 dataset and Recall for the Charades-STA dataset, is more pronounced. This suggests that the performance gain of the VTG-LLM primarily stems from more accurate timestamp location. Additionally, performance on other metrics like CIDEr and SODA_c also improves.
Table 1: Zero-shot performance of algorithms over various tasks. We evaluated the performance of VTG-LLM using the Youcook2, Charades-STA, and QVHighlights datasets. We highlight the best results for each block using bold font. The Valley, VideoChat-Embed, Video-ChatGPT, and Video-LLaMA results are elaborated from previous studies (Ren et al., 2023; Huang et al., 2023; Qian et al., 2024). The results with gray color indicates unfair comparison (13B). The VTG-LLM samples 96 frames during training and testing stages.
Model Youcook2 Charades-STA QVHighlights
[…]
Traditional Video LLMs
Valley (7B) 0.1 0.0 1.5 4.7 1.6 10.9 15.2
VideoChat (7B) 0.2 0.6 3.4 3.2 1.4 13.1 18.1
Video-LLaMA (7B) 0.0 0.0 0.1 2.7 1.2 11.3 15.6
Video-ChatGPT (7B) 7.7 1.7 3.8
Temporal Grounding Video LLMs
TimeChat (7B) 1.2 3.4 12.6 32.2 13.4 14.5 23.9
VTimeLLM (7B) 27.5 11.4
VTimeLLM (13B) 34.3 14.7
Momentor (7B) 26.6 11.6 7.6
HawkEye (7B) 31.4 14.5
VTG-LLM (7B) 1.5 5.0 17.5 33.8 15.7 16.5 33.5
Quality Analysis.
In Figure 3, we compare the outputs of VTG-LLM and TimeChat in a zero-shot manner on the Youcook2 dataset. We observe that VTG-LLM’s outputs are more closely aligned with human intention. For instance, there are no repeated sentences with the same meaning.
Figure 3: Zero-shot case study. We compare the model outputs of VTG-LLM and TimeChat (Ren et al., 2023) on Youcook2 dataset in the zero-shot manner. The colored texts indicating repeated captions. More case studies can be found in Appendix C.
4.3 Ablation Studies
In this section, we aim to show that all the designed components are necessary for VTG-LLM to achieve the best results.
Table 2: Ablation studies on VTG-LLM. We perform ablation studies on the three components of the algorithm. For the Sequence-Time Embedding section, "SE Only" refers to using only sequence embedding. "TE Random Initialize" represents the random initialization of time embedding rather than the zero initialization we recommend. In the Time Token component, "No Time Token" signifies not using the time token, while "Time Token not Formatted" implies not formatting timestamps to the same length. "Token Embedding not Initialized" indicates randomly initialize the token embedding and LLM prediction head for time tokens. Lastly, for the Token Compression aspect, we compare slot-based token compression to several straightforward baselines. We highlight the best results of ablation studies using underline.
Model Youcook2 Charades-STA QVHighlights
[…]
Ablation Studies on STE
SE Only 1.4 4.4 17.2 33.6 14.8 16.4 32.8
SE + TE (Random Initialize) 1.3 4.2 16.8 21.4 10.5 14.2 22.5
Ablation Studies on Time Token
No Time Token 1.7 5.5 17.4 29.5 13.6 16.2 33.9
Time Token not Formatted 1.5 5.1 16.8 27.0 12.4 15.0 28.8
Token Embedding not Initialized 1.4 4.8 16.1 15.6 6.5 14.0 21.9
Ablation Studies on Token Compression
Entropy Sampling 1.3 4.2 16.8 21.1 9.6 14.0 20.1
Diverse Sampling 1.3 4.4 16.8 22.4 10.5 14.2 22.7
Cross Attention 1.3 4.2 17.6 19.9 8.8 13.7 21.0
Original VTG-LLM 1.5 5.0 17.5 33.8 15.7 16.5 33.5
Ablation studies on types of token compression methods.
In the "Ablation Studies on Token Compression" section of Table 2, we present the performance of VTG-LLM using different token compression methods and include some naive baselines. The "Entropy Sampling" method involves sampling the number of tokens to 256 by selecting tokens with the maximum entropy, where the entropy is estimated using k-nearest neighbor distances (Kozachenko & Leonenko, 1987). The "Diverse Sampling" method involves sampling tokens to 256 using k-means++ (Arthur et al., 2007) to select tokens with the maximum distance. The "Cross Attention" method involves using a cross-attention layer to compress the number of tokens. The results indicate that: (1) The slot-based sampling achieves better performance among all compression methods; (2) Sampling-based methods perform poor compared to slot-based token compression method. We conjecture that this is because the sampling methods drop too much necessary information; (3) The cross-attention method performs even worse than sampling-based methods, which might be attributed to the attention mechanism requiring a large amount of data to perform well.
Ablation studies on sequence-time embedding.
In the "Ablation Studies on STE" section of Table 2, we show the performance of VTG-LLM when only using sequence embedding ("SE Only") and using sequence-time embedding but randomly initializing the weights of time-embedding ("TE Random Initialize"). The results indicate that: (1) the time-embedding is necessary for VTG-LLM to achieve the best result. Without the time-embedding, the performance of VTG-LLM on all datasets decreases; (2) Randomly initializing the time embedding significantly hurts the performance of VTG-LLM, making its performance worse than "SE Only".
Ablation studies on time-token.
In the "Ablation Studies on Time Token" section of Table 2, we show the performance of VTG-LLM when not using special time tokens ("No Time Token"), when not formatting the length of timestamps to six time tokens ("Time Token not Formatted"), and when do not initialize the token embedding and LLM head for time tokens ("Token Embedding not Initialized"). The results indicate that
[…]
  • We found that the "No Time Token" setting tends to predict fewer timestamps on QVHighlights dataset, which may lead to better HIT@1 performance but a worse mAP score.
  • Formatting the timestamps using the same number of time tokens significantly enhances the model performance ("Time Token not Formatted"). We believe that this is because the formatted setting simplifies the task, making the models needless to identify the number of tokens in timestamps.
  • Initialize the weights of token embedding and LLM prediction head is essential for time token to work well.
In summary, we recommend using special time tokens to achieve significant improvements in the accuracy of locating timestamps, which is the primary goal of this paper. Moreover, investigating the trade-off between caption quality and timestamp location accuracy in future work would be interesting.
Table 3: Fine-tuned performance of algorithms over various tasks. We fine-tune the Algorithm on Youcook2, Charades-STA, and QVHighlights datasets with a batch size of 16. We use gray color for traditional non-generalist models and task-specific models. Vid2Seq (Yang et al., 2023) fully training the LLM models and using 1B video data for pretraining.
Model Youcook2 Charades-STA QVHighlights
[…]
Non-Generative Models
InternVideo2-6B (Wang et al., 2024a) 70.0 49.0
VDI (Luo et al., 2023a) 52.3 31.4
Moment-DETR (Lei et al., 2021) 55.7 34.2 37.4 60.2
HL-CLIP (Han et al., 2024) 41.9 70.6
Task-Specific Models
Vid2Seq (Yang et al., 2023) 7.9 47.1 27.3
Vid2Seq (Visual Only) (Yang et al., 2023) 5.7 25.3 23.5
HawkEye 58.3 28.8
Generalist Models
TimeChat 3.4 11.0 19.5 46.7 23.7 21.7 37.9
VTG-LLM 3.6 13.4 20.6 57.2 33.4 24.1 41.3
Fine-tuned model performance of VTG-LLM.
In Table 3, we fine-tune the VTG-LLM models on Youcook2, Charades-STA, and QVHighlights datasets. From the results, we make the following observations: (1) The performance of VTG-LLM is significantly better than other generalist models across all three tasks; (2) The performance of VTG-LLM on the moment retrieval task (Charades-STA) is comparable to task-specific models like HawkEye, as well as non-generative models like VDI and Moment-DETR. (3) The task-specific model Vid2Seq (Yang et al., 2023) achieves superior performance on the Youcook2 dataset by fully training the LLM models and using 1B video data for pretraining. In contrast, we only fine-tune the LLM using LoRA and employ 120K instruction tuning data.
5 Conclusion, Limitation, and Future Works
In this paper, we introduce VTG-IT-120K, a high-quality instruction tuning dataset that covers four video temporal grounding (VTG) tasks, including moment retrieval, dense video captioning, video highlight detection, and video summarization. Furthermore, we propose the VTG-LLM, an improved video LLM model for integrating knowledge about timestamps, enhancing the model performance on VTG tasks. Extensive numerical results demonstrate the superior zero-shot performance of VTG-LLM over state-of-the-art video LLM models on Charades-STA, QVHighlights, and Youcook2 datasets, with each proposed component contributing to individual performance gains.
However, there are still some constraints with the VTG-LLM that warrant further research. For instance, the VTG-LLM’s capacity for dense video captioning and highlight detection remains subpar compared to non-generative or task-specific models. Therefore, exploring more advanced methods to rectify this issue is a worthwhile endeavor. Additionally, while this study primarily focuses on the accuracy of timestamp location, it would also be beneficial to investigate techniques to enhance other aspects, such as the precision of salient scores and the quality of captions. Moreover, we did not incorporate the audio elements of the videos in this study. Integrating these in the future could potentially enhance the overall performance.
Appendix A Experiments
A.1 Detailed Experiment Settings
In this section, we present the detailed experimental settings used in our study. We employ ViT-G/14 from EVA-CLIP (Sun et al., 2023) and Qformer from InstructBLIP (Dai et al., 2024) as the visual backbone. The number of queries for Qformer is set to 32. The vision projection layer and the weights of the sequence embedding are initialized using the Video-LLaMA checkpoint (Zhang et al., 2023). To address the mismatched size of the sequence embedding, we use interpolation to extend the size from 32 to 96. The time-embedding (8192) is initialized with zeros. We set the number of slots to 256.
For the language model backbone, we utilize LLaMA-2-7B (Touvron et al., 2023). The LoRA training is conducted on "q_proj", "k_proj", "v_proj", and "o_proj". Special time tokens are added to the LLaMA tokenizer, and we set the prediction head of LLM and the weights of token embedding to be trainable. The maximum text length is set to 2048.
The models are trained on the VTG-IT-120k and a randomly sampled subset (97k) from the Valley dataset (Luo et al., 2023b). We randomly sample 96 frames for each video. Training is carried out using 16 ATN 910B GPUs with a batch size of 64 and takes approximately 40 hours to complete. The training employs DDP and requires about 30GB of GPU storage for each GPU. Fine-tuning also uses 16 GPUs and sets the batch size to 16. We fine-tune the VTG-LLM model for 10 epochs on Charades dataset, 16 epochs for Youcook2 dataset, and 20 epochs for QVHighlights dataset. We set the initial learning rate to 3e-5 and train the models for 10 epochs by default. The weight decay is set to 0.05, and the warm-up steps use $0.6E$, where $E$ represents the number of iterations for each epoch.
Appendix B Data Annotation Examples
We present annotation examples for moment retrieval (Figure 4), dense video captioning (Figure 5), video highlight detection (Figure 6), and video summarization (Figure 7). It should be noted that in the VTG-IT-VHD dataset, only clips with high scores are annotated. This leads to a relatively consistent score distribution among the annotated clips. As a result, the VTG-IT-VHD dataset assists models in locating relevant video clips; however, the significant scores may not be accurately predicted in a zero-shot manner due to differing scoring criteria. This issue can be addressed by fine-tuning the models on target datasets, which helps the models understand the specific scoring criteria.
Figure 4: Annotation example of moment retrieval task.
Figure 5: Annotation example of dense video captioning task.
Figure 6: Annotation example of dense video highlight detection task. It should be noted that in the VTG-IT-VHD dataset, only clips with high scores are annotated. This leads to a relatively consistent score distribution among the annotated clips.
Figure 7: Annotation example of dense video summarization task.
Appendix C Additional Case Studies
We show the additional zero-shot case studies on dense video captioning (Figure 8), moment retrieval (Figure 9), and video highlight detection (Figure 10) tasks.
Figure 8: Zero-shot case study of dense video captioning task.
Figure 9: Zero-shot case study of moment retrieval task.
Figure 10: Zero-shot case study of video highlight detection task.